OpenStudy (asad786):

What is the solution to the open sentence? 4y – 3 = y + 6 A. 1 B. 2 C. 3 D. 9

OpenStudy (anonymous):

A?

OpenStudy (kitkats):

C

OpenStudy (kitkats):

4y - 3 = y+6 Subtract y from both sides 3y - 3 = 6 Add 3 from both sides 3y = 9 Divide 3 from both sides y = 3